OpenStudy (anonymous):

x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (goldphenoix):

Do the distributive property. 3 * x = ? 3 * 1 = ?

OpenStudy (anonymous):

Im still not understanding

OpenStudy (goldphenoix):

What does 3 * 1 give you?

OpenStudy (anonymous):

3

OpenStudy (goldphenoix):

Good. Now what does 3 * x give you?

OpenStudy (anonymous):

3x

OpenStudy (goldphenoix):

Good. 3(x + 1) would become 3x + 3. Now do the same for −2(x − 1) -2 * x = ? -2 * -1= ?

OpenStudy (anonymous):

so it would be 3x +3 = -2x - -4

OpenStudy (goldphenoix):

Hmm. -2 * -1 = ?